OpenStudy (anonymous):

is the mean of 42,50,55 49?

OpenStudy (anonymous):

Sorry, messed that up. Yes, 42+50+55=147 147/3=49

OpenStudy (anonymous):

Thanks and its fine :)

OpenStudy (anonymous):

The mean is the average. So 42+50+55=147 the you take the amount of numbers and divide by the answer so 147/3=49 49 is your answer. please mark as helpful :)

OpenStudy (anonymous):

oh opps i didnt realize that someone already answered your question sorry